HUBBLE SPACE TELESCOPE DAILY REPORT #1155 PERIOD COVERED: 0700 EST Daily Status Report as of 159/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 5528 (FOS Cycle 4 Focus, X-Pitch, Y-Pitch) The Faint Object Spectrograph (Blue Side) was used to make both a TALED and a WAVE calibration observation. No external sources were used. The observations were received and displayed. The observations were reported to look okay. 1.2 Completed Two Sets of WF/PC-2 5655 (WF/PC-2 Cycle Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make both dark and visual flats. Three images from the first set were received and displayed. One was partially saturated, one was underexposed, and the other was okay. Data for the second set of observations still remains on Science Tape Recorder #3, and will be analyzed later. No anomalies were noted. 1.3 Completed FOS/BL 5579 (Cycle 4 Test: To Determine FOS Polarimetry Mode) The Faint Object Spectrograph (Blue Side) was used to make calibration observations of star BD+28D4211 to study its polarization capabilities. Three observations were executed as scheduled. The peakups received maximum counts in groups 4 (6X2), 9 (6X2), and 2 (5X5) respectively. No anomalies were noted. 1.4 Completed WF/PC-2 5092 (GTO Low Latitude Parallel Program: Cycle 4) The WF/PC-2 was used to make observations taken in parallel with the prime science observations. This is part of a survey to take pictures of a large number of fields within 30 degrees of the galactic plane. The observations were executed as scheduled, and the science data remains on Science Tape Recorder #1. It will be analyzed later. 1.5 Completed WF/PC-2 5563 (WF/PC-2 Cycle 4 Photometry Calibration Monitor 1: UV/OPT STD.) The WF/PC-2 was used to make photometric calibration images of ultraviolet standard star GRW+70D582 and monitor the photometric stability and quantum efficiency of the WF/PC-2. The observations completed without apparent anomaly. 1.6 Completed WF/PC-2 5476 (Continuum Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C272.0 as part of a continuum survey and to study the galaxy's morphology. The observations completed without apparent anomaly. 1.7 Completed WF/PC-2 5643 (WF/PC-2 Cycle Calibration: Partially Rotated Filters) The WF/PC-2 was used to make an internal flat observations. The observations was executed as scheduled, and the data remains on Science Tape Recorder #1. It will be analyzed later. 1.8 Completed FOS/BL 5438 (Io's Atmosphere: UV Spectroscopy of Neutral and Molecular Components) The Faint Object Spectrograph (Blue side) was used to make far ultraviolet spectrographic observations of Jupiter's moon Io. The baseline guide-star acquisition was achieved in fine lock, and the decision observation proceeded on a tracking slew with over 1.5 million counts in the overlight sum. After breaking down the data file into five groups, the peak signal was seen in the first group, which would imply that the target was at or near the edge of the aperture. The spectrum looked as expected. A corrective SIP maneuver was successfully uplinked to the spacecraft. The subsequent observation occurred during LOS. 1.9 Completed WF/PC-2 5099 (Imaging and Spectroscopy of a Complete Sample of Bright Nearby Quasars) The WF/PC-2 was used to image nearby quasar PG1301-012. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled :11 Successful :9 HSTAR 4800 at 159/0626Z, FHST#2 failed its error box check during LOS. At AOS there was an indication that the maneuvers to null the errors observed by FHST #2 had not occurred. 2.3 Significant HSTARs: None 2.4 Operations Notes: During the last day, HST operations have been conducted uneventfully except for the FHST update failures mentioned above. 3.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-